I tryed solving this problem and im just not getting the right answer, can you tell me what i am doing wrong? :
3/4(x-8)=3(x+6) This is the problem
where can i send my work?

Answer by jim_thompson5910(35256)

You can post the work along with the problem (if possible)

3/4(x-8)=3(x+6)

3(x-8)=4*3(x+6)

3(x-8)=12(x+6)

3x-24=12x+72

3x=12x+72+24

3x=12x+96

3x-12x=96

-9x=96

x=96/(-9)

x = -32/3